• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:08:45 on January 22, 1999

    Entry number 9665

    keyword=Beam Time Accounting
    Hall A

    1/22/1999 0000-0800
    Experiment #: E94-010
    Shift Leader: Feng Xiong
    Shift Workers: Z.Chai, , P. Djawotho

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    0 0 0 0 25 60 60 0 2.41666666666667
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 60 1
    Beam Not Available
    or Unacceptable
    60 60 60 60 35 0 0 0 4.58333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    60 60 60 60 60 60 60 0 7
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 60 1
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: From 12:30 AM to 4:35AM, we were waiting for the energy stabilization. BCM didn't work from 4:35 AM to 6:00AM; EPICS system for laser and RTD readout failed until now. We finished the raster setup. We finished empty cell yield for five different position, and found the lowest background is from x=0, y=0. Changed the kinematic settings for electron arm, theta=35.5, p=0.667 GeV/c.

    Data Entry by: F Xiong